OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. -5, -25, -125, -625, ...

OpenStudy (igigighjkl):

do you want the awnser?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

if you know it

OpenStudy (igigighjkl):

I do, 3125, 15625, 78125. Was this helpful?

OpenStudy (anonymous):

i need the rule

OpenStudy (igigighjkl):

x 5.

OpenStudy (igigighjkl):

Was this helpful?

OpenStudy (igigighjkl):

Please reply.

OpenStudy (anonymous):

that isnt one of my answer choices

OpenStudy (igigighjkl):

really? what are they.

OpenStudy (anonymous):

an = -5 • 5^n an = 5 • -5^n + 1 an = 5 • -5^n an = -5 • 5^n - 1 I think its D

OpenStudy (anonymous):

but i am not sure

OpenStudy (igigighjkl):

what one second.

OpenStudy (igigighjkl):

-5x5^n

OpenStudy (igigighjkl):

was that helpful?

OpenStudy (igigighjkl):

please reply.

OpenStudy (igigighjkl):

hello?

OpenStudy (igigighjkl):

@alejandranunez

OpenStudy (anonymous):

thank you

OpenStudy (igigighjkl):

your welcome at anytime.